OpenStudy (anonymous):

simplify (2√5+3√7)^2 Justify your answer I posted this with no answer earlier. I hope i can get one now

OpenStudy (tyteen4a03):

Remember that \((a + b)^2 = a^2 + 2ab + b^2\).

OpenStudy (luigi0210):

It's the same as: \[(2\sqrt{5} +3\sqrt{7})*(2\sqrt{5} +3\sqrt{7})\] Just multiply it out

OpenStudy (anonymous):

Okay, thanks

OpenStudy (luigi0210):

It should be: \[4(5)+6(\sqrt{35})+6(\sqrt{35})+9(7)\] Now just simplify

OpenStudy (anonymous):

20 + 36 + 36 + 63 = 155. Answer?

OpenStudy (luigi0210):

Actually it be \[83+12\sqrt{35}\]

OpenStudy (anonymous):

oh.. so then whats the answer? Or is that the answer?

OpenStudy (luigi0210):

That's the answer.

OpenStudy (tyteen4a03):

@HatSimulator How did you get to 20 + 36 + 36 + 63? Also, @Luigi0210 's step can be further simplified using my identity.

OpenStudy (luigi0210):

Wait, you sure?

OpenStudy (tyteen4a03):

@Luigi0210 Yes. Do you need proof?

OpenStudy (anonymous):

I'd like proof

OpenStudy (tyteen4a03):

(a + b)^2 = (a + b)(a + b) = a^2 + ab + ab + b^2 = a^2 + 2ab + b^2